# Better Late Than Never: Formulating and Benchmarking Recommendation Editing

### v1

@@ A.5 Editing Method
Finetuning-based Approaches: FT (Fine-Tuning) Sinitsin et al. [2020] and EFT (Editing Fine-Tuning). These methods edit the model using a fine-tuning manner based on explicit editing data. For example, the FT method modifies the recommendation model by changing its original parameters, requiring the entire graph’s information for graph-based collaborative filtering models. In contrast, EFT modifies the final user and item embeddings through a model that generates these embeddings, without requiring detailed knowledge of the model’s architecture and training data.
@@
More specifically, the FT method transmits the gradient of the edit loss to the original parameters of the collaborative filtering model and then utilizes the Adam optimizer to modify the model parameters for editing. The EFT method pre-fetches the user and item embeddings generated by the original collaborative filtering model. After parameterizing these embeddings, the gradient of the edit loss is transmitted solely to these embeddings. Editing is then performed by modifying these embeddings. Algorithm 1 describes the process of editing using the E-BPR Loss with EFT.
Regularization-based Approaches: LWF (Learning Without Forgetting) Li and Hoiem [2017], L2 (Online L2Reg) Lin et al. [2022], SRIU (Sample Reweighting Incremental Update) Peng et al. [2021]. These methods also edit models by fine-tuning explicit editing data and using the original model for regularization during the editing process. LWF normalizes the output on the original model using editing data, and L2 utilizes the parameters of the original model for regularization. SRIU contends that overfitting editing data with low confidence in the original model can lead the edited model to deviate from the original model’s patterns. It adopts a re-weighting approach that assigns lower weights to data with low confidence.
@@
More specifically, LWF and L2 are extensions of the FT method used for editing with E-BPR loss, where an additional regularization loss is introduced to prevent knowledge forgetting. LWF regularizes the editing process by using the MSE loss between the scores of explicit edit pairs from the original and edited models. Conversely, L2 directly employs the MSE loss between the parameters of the original model and the edited model for regularization. The formulations of their edit losses are as follows:
$\mathcal{L}=\mathcal{L}_{e\_bpr}+\lambda\mathcal{L}_{reg}.$ (8)
@@
SRIU normalizes the output scores of the original model on explicit edit pairs using mean and variance normalization. Then, it computes the weight for each explicit edit pair through Softmax. For the E-BPR loss generated for each triplet (u, $i_{pos}$, $i_{neg}$), the corresponding loss is multiplied by the weight of its associated explicit edit pair.
@@
Replay-based Approaches: RSR (Random Sampling with a Reservoir) Vitter [1985] and SPMF (Stream-centered Probabilistic Matrix Factorization model) Wang et al. [2018]. This category method balances the efficiency and effectiveness of model editing by sampling a small number of training examples. The model is edited by fine-tuning based on both the sampled data and the explicit editing data. RSR adopts random sampling, while SPMF uses the original model’s prediction scores on training samples to perform weighted sampling.RSR selects $N$ sets of positive feedback at random from historical training data. Conversely, SPMF calculates weights based on the output scores of the original model on the training data and performs weighted sampling to select $N$ sets of positive and negative feedback. Editing is then conducted using the BCE loss.
@@
Optimization-based Approaches: SML (Sequential Meta-Learning) Zhang et al. [2020]. SML edits the model by training a Transfer module to generate the new parameters after model editing. Due to the constraints of our task, we cannot train SML on sequential data as originally designed for SML. Instead, we simulate editing data by randomly sampling the recommendation results of the original model multiple times. This simulated data is then used to train the Transfer module for the editing process. The original SML combines the parameters of the original model with those trained on incremental data as inputs for the parameter transformer. However, in our task, due to the small number of explicit edit pairs, which only contain negative feedback, it is impractical to train a reasonable recommendation model. Therefore, we modify the parameters trained on incremental data to be those obtained from fine-tuning the original model on explicit edit pairs for one round.
@@
CF with Negative Feedback: SiReN (Sign-aware Recommender system based on GNN models) Seo et al. [2022] and SiGRec (Signed Graph Neural Network Recommendation model) Huang et al. [2023]. These methods handle both positive and negative feedback during collaborative filtering model training. In this paper, we adopt these approaches to handle negative feedback for model editing purposes.SiReN requires an additional MLP module to process negative feedback. After obtaining the negative feedback embeddings, they are merged with the positive feedback embeddings using an attention mechanism. The MLP module and attention parameters are pre-trained by simulating real negative feedback through random sampling of negative feedback instances.SiGRec aggregates negative embeddings through message passing on a negative graph, which are then concatenated with the user and item embeddings obtained from the original collaborative filtering model. This process results in new user and item embeddings. SiGRec uses the Sign Cosine loss function proposed in the original paper for editing.
@@
Editing Approach: EGNN (Editable Graph Neural Networks) Liu et al. [2023] and BiEGNN proposed by ourselves. EGNN enhances editing capabilities through an additional MLP (Multi-Layer Perceptron) module. Given the lack of node attributes in collaborative filtering tasks, we use the representations generated by the original model as inputs for the editing module. The existence of two types of nodes (i.e., user and item) utilizing the same editing module may create bias. To tackle this issue, we propose an improved version of EGNN, named BiEGNN, which trains distinct editing modules for users and items respectively to conduct editing. EGNN and BiEGNN employ an MLP as the editing module, with the user and item embeddings obtained from the original collaborative filtering model serving as inputs to the editing module. Before editing, it is necessary to pre-train the editing module. This pre-training is conducted by minimizing the score differences between the original model and the modified model on the training data, using the MSE loss.
A.6 Editing Method Parameter Setting
@@
A.7 Hardware and software configuration
All experiments are executed on a server with 755GB main memory, and two AMD EPYC 7763 CPUs. All experiments are done with a single NVIDIA RTX 4090 (24GB). The software and package version is specified in Table 3.
@@ A.7 Hardware and software configuration
Table 3: Package configurations of our experiments.
Appendix B More Experimental Results
B.1 Recommendation Models Performance Before Editing
B.2 Editing Effectiveness
Table 5 and Table 6 display a detailed comparison of various methods across different datasets and backbones, showcasing the EC, EP, Recall, and NDCG metrics. We can find that the results are consistent with the main analysis in Section 4.1.
B.3 Editing Efficiency
Figure 5: Editing efficiency.
@@ B.3 Editing Efficiency
Table 4: Recommendation models performance before editing.
Table 5: Experimental results (Mean ± Std.) of EC and EP metrics for all methods.
Epinions MF LightGCN XSimGCL
Method EC EP EC EP EC EP
FT 0.6485±0.0236 0.8224±0.0365 0.4580±0.0183 0.7879±0.0531 0.5932±0.0172 0.7276±0.0504
LWF 0.6554±0.0284 0.8080±0.0429 0.4494±0.0174 0.8094±0.0465 0.5949±0.0223 0.7250±0.0496
L2 0.6610±0.0247 0.8001±0.0354 0.4895±0.0117 0.8702±0.0183 0.5054±0.0213 0.9481±0.0141
SRIU 0.6564±0.0268 0.7997±0.0457 0.4498±0.0134 0.7612±0.0556 0.6068±0.0208 0.7073±0.0511
RSR 0.7936±0.0117 0.8823±0.0338 0.5800±0.0190 0.5236±0.0876 0.5853±0.0147 0.7376±0.0488
SPMF 0.7976±0.0123 0.8919±0.0243 0.5768±0.0179 0.5338±0.0916 0.5932±0.0159 0.7411±0.0495
SML 0.8230±0.0001 0.0010±0.0000 0.6907±0.0001 0.1216±0.0000 0.6983±0.0000 0.1036±0.0000
SiReN 0.5359±0.0331 0.6265±0.1316 0.6135±0.0771 0.2930±0.1859 0.5406±0.0491 0.4914±0.1653
SiGRec 0.0007±0.0003 0.9342±0.0297 0.6682±0.0612 0.2772±0.1175 0.6754±0.0345 0.4770±0.1080
EGNN 0.5629±0.0256 0.5293±0.0943 0.5866±0.0265 0.5082±0.0687 0.5455±0.0557 0.5361±0.1354
BiEGNN 0.6495±0.0275 0.8182±0.0387 0.4470±0.0239 0.8318±0.0194 0.5968±0.0199 0.7178±0.0593
EFT 0.6502±0.0390 0.8193±0.0482 0.6920±0.0200 0.8296±0.0444 0.5813±0.0368 0.8906±0.0271
KuaiRand MF LightGCN XSimGCL
Method EC EP EC EP EC EP
FT 0.5355±0.0333 0.7172±0.0131 0.4421±0.0135 0.7882±0.0201 0.4265±0.0166 0.8300±0.0331
LWF 0.5350±0.0342 0.7189±0.0113 0.4428±0.0107 0.7806±0.0387 0.4287±0.0147 0.8296±0.0366
L2 0.5356±0.0335 0.7177±0.0129 0.4129±0.0255 0.9248±0.0235 0.4490±0.0167 0.7057±0.0192
SRIU 0.5350±0.0326 0.7174±0.0145 0.4401±0.0186 0.7467±0.0432 0.4314±0.0210 0.7908±0.0435
RSR 0.6493±0.0362 0.8702±0.0327 0.4831±0.0169 0.5132±0.0592 0.4367±0.0184 0.8237±0.0370
SPMF 0.6786±0.0377 0.8836±0.0323 0.4915±0.0161 0.4865±0.0656 0.4361±0.0187 0.8253±0.0382
SML 0.6486±0.0003 0.0564±0.0008 0.6340±0.0000 0.1871±0.0000 0.6279±0.0000 0.1756±0.0000
SiReN 0.4834±0.0316 0.5346±0.1299 0.5772±0.0818 0.3565±0.1685 0.6562±0.0441 0.2642±0.1340
SiGRec 0.0006±0.0002 0.9247±0.0271 0.6114±0.0476 0.2349±0.0772 0.5375±0.0456 0.4804±0.0953
EGNN 0.4682±0.0357 0.6083±0.1040 0.5236±0.0369 0.4398±0.0870 0.5119±0.0288 0.5119±0.0753
BiEGNN 0.5363±0.0334 0.7148±0.0165 0.4412±0.0097 0.7966±0.0316 0.4272±0.0165 0.8148±0.0361
EFT 0.5358±0.0338 0.7183±0.0142 0.5897±0.0212 0.8286±0.0275 0.5192±0.0237 0.8453±0.0109
QB-video MF LightGCN XSimGCL
Method EC EP EC EP EC EP
FT 0.7370±0.0209 0.7779±0.0486 0.4829±0.0319 0.8023±0.0474 0.5028±0.0267 0.7803±0.0311
LWF 0.7390±0.0217 0.7657±0.0494 0.4773±0.0351 0.8263±0.0554 0.4910±0.0331 0.8085±0.0612
L2 0.7070±0.0131 0.8307±0.0282 0.4535±0.0187 0.9669±0.0070 0.4808±0.0270 0.8451±0.0696
SRIU 0.7406±0.0255 0.7620±0.0506 0.4651±0.0264 0.8139±0.0357 0.5448±0.0359 0.6865±0.0428
RSR 0.7367±0.0502 0.8794±0.0346 0.5750±0.0212 0.4657±0.0551 0.5207±0.0393 0.7081±0.0550
SPMF 0.8050±0.0152 0.8097±0.0181 0.5748±0.0295 0.4742±0.0846 0.5176±0.0312 0.7127±0.0341
SML 0.9000±0.0000 0.0001±0.0000 0.6962±0.0000 0.1376±0.0000 0.6870±0.0000 0.1105±0.0000
SiReN 0.5617±0.0449 0.4192±0.1517 0.6531±0.1108 0.3363±0.2369 0.6571±0.0849 0.3194±0.2502
SiGRec 0.0010±0.0004 0.9007±0.0464 0.6954±0.0604 0.3006±0.0811 0.5894±0.0372 0.5371±0.0376
EGNN 0.5243±0.0434 0.6291±0.0558 0.5722±0.0309 0.5001±0.0757 0.5628±0.0397 0.5231±0.0386
BiEGNN 0.7409±0.0231 0.7691±0.0300 0.4849±0.0224 0.7942±0.0289 0.4988±0.0383 0.7794±0.0843
EFT 0.7394±0.0245 0.7670±0.0429 0.6922±0.0156 0.8920±0.0255 0.5733±0.0195 0.9224±0.0043
Table 6: Experimental results (Mean ± Std.) of Recall and NDCG metrics for all methods.
Epinions MF LightGCN XSimGCL
Method Recall NDCG Recall NDCG Recall NDCG
FT 0.1312±0.0008 0.0607±0.0003 0.1564±0.0010 0.0741±0.0006 0.1692±0.0019 0.0835±0.0007
LWF 0.1308±0.0010 0.0605±0.0005 0.1569±0.0010 0.0743±0.0005 0.1688±0.0026 0.0834±0.0010
L2 0.1306±0.0006 0.0605±0.0003 0.1540±0.0006 0.0734±0.0004 0.1737±0.0004 0.0851±0.0001
SRIU 0.1306±0.0012 0.0604±0.0006 0.1570±0.0013 0.0741±0.0006 0.1680±0.0019 0.0829±0.0010
RSR 0.1301±0.0014 0.0595±0.0009 0.1367±0.0071 0.0653±0.0033 0.1691±0.0019 0.0835±0.0009
SPMF 0.1300±0.0011 0.0596±0.0007 0.1373±0.0070 0.0657±0.0033 0.1696±0.0019 0.0836±0.0010
SML 0.0021±0.0000 0.0008±0.0000 0.0776±0.0000 0.0355±0.0000 0.0779±0.0000 0.0356±0.0000
SiReN 0.1228±0.0068 0.0566±0.0036 0.0945±0.0377 0.0445±0.0182 0.1479±0.0203 0.0715±0.0100
SiGRec 0.1304±0.0015 0.0602±0.0008 0.1036±0.0230 0.0490±0.0109 0.1484±0.0120 0.0738±0.0057
EGNN 0.1181±0.0059 0.0548±0.0028 0.1389±0.0062 0.0662±0.0027 0.1556±0.0138 0.0761±0.0075
BiEGNN 0.1311±0.0008 0.0607±0.0004 0.1573±0.0008 0.0745±0.0003 0.1685±0.0028 0.0832±0.0013
EFT 0.1310±0.0008 0.0606±0.0003 0.1563±0.0011 0.0741±0.0005 0.1726±0.0006 0.0847±0.0002
KuaiRand MF LightGCN XSimGCL
Method Recall NDCG Recall NDCG Recall NDCG
FT 0.1764±0.0015 0.0656±0.0006 0.2329±0.0013 0.0875±0.0006 0.2494±0.0018 0.0956±0.0006
LWF 0.1766±0.0018 0.0657±0.0006 0.2328±0.0012 0.0877±0.0008 0.2493±0.0016 0.0956±0.0006
L2 0.1759±0.0019 0.0654±0.0008 0.2325±0.0016 0.0880±0.0004 0.2458±0.0023 0.0941±0.0008
SRIU 0.1764±0.0013 0.0656±0.0005 0.2326±0.0014 0.0875±0.0008 0.2492±0.0023 0.0952±0.0006
RSR 0.1808±0.0013 0.0670±0.0006 0.2033±0.0079 0.0768±0.0029 0.2486±0.0020 0.0954±0.0006
SPMF 0.1804±0.0012 0.0668±0.0004 0.1996±0.0084 0.0754±0.0033 0.2489±0.0018 0.0956±0.0007
SML 0.0591±0.0008 0.0189±0.0002 0.1649±0.0000 0.0613±0.0000 0.1646±0.0000 0.0611±0.0000
SiReN 0.1588±0.0129 0.0587±0.0050 0.1700±0.0492 0.0625±0.0198 0.1532±0.0482 0.0574±0.0188
SiGRec 0.1793±0.0012 0.0660±0.0007 0.1536±0.0219 0.0563±0.0091 0.2116±0.0157 0.0815±0.0063
EGNN 0.1695±0.0084 0.0631±0.0038 0.1947±0.0164 0.0736±0.0061 0.2234±0.0122 0.0857±0.0041
BiEGNN 0.1762±0.0017 0.0657±0.0006 0.2326±0.0015 0.0876±0.0009 0.2481±0.0019 0.0952±0.0005
EFT 0.1768±0.0016 0.0657±0.0005 0.2315±0.0013 0.0876±0.0004 0.2473±0.0004 0.0951±0.0002
QB-video MF LightGCN XSimGCL
Method Recall NDCG Recall NDCG Recall NDCG
FT 0.2893±0.0033 0.1748±0.0025 0.3182±0.0010 0.1965±0.0012 0.3462±0.0014 0.2148±0.0010
LWF 0.2888±0.0032 0.1747±0.0023 0.3185±0.0009 0.1968±0.0008 0.3477±0.0027 0.2159±0.0024
L2 0.2915±0.0019 0.1761±0.0015 0.3193±0.0002 0.1972±0.0002 0.3487±0.0019 0.2169±0.0014
SRIU 0.2887±0.0028 0.1744±0.0020 0.3185±0.0007 0.1970±0.0006 0.3383±0.0031 0.2100±0.0024
RSR 0.2910±0.0021 0.1731±0.0032 0.2688±0.0114 0.1605±0.0075 0.3397±0.0048 0.2109±0.0036
SPMF 0.2855±0.0017 0.1660±0.0024 0.2691±0.0168 0.1611±0.0112 0.3404±0.0031 0.2114±0.0021
SML 0.0003±0.0000 0.0002±0.0000 0.1686±0.0000 0.1015±0.0000 0.1648±0.0000 0.1000±0.0000
SiReN 0.2258±0.0464 0.1317±0.0311 0.1926±0.1093 0.1166±0.0678 0.2100±0.1086 0.1290±0.0690
SiGRec 0.2882±0.0034 0.1727±0.0027 0.2377±0.0266 0.1398±0.0173 0.3209±0.0099 0.1981±0.0080
EGNN 0.2767±0.0073 0.1652±0.0064 0.2814±0.0146 0.1713±0.0103 0.3062±0.0116 0.1890±0.0085
BiEGNN 0.2893±0.0017 0.1751±0.0011 0.3172±0.0021 0.1957±0.0021 0.3456±0.0038 0.2144±0.0028
EFT 0.2890±0.0024 0.1746±0.0018 0.3186±0.0005 0.1967±0.0006 0.3502±0.0003 0.2176±0.0002

### v2

@@ A.3. Editing Method
FT: The FT method transmits the gradient of the edit loss to the original parameters of the collaborative filtering model and then utilizes the Adam optimizer to modify the model parameters for editing.
EFT: The EFT method pre-fetches the user and item embeddings generated by the original collaborative filtering model. After parameterizing these embeddings, the gradient of the edit loss is transmitted solely to these embeddings. Editing is then performed by modifying these embeddings. Algorithm 1 describes the process of editing using the E-BPR Loss with EFT.
@@
LWF and L2: LWF and L2 are extensions of the FT method used for editing with E-BPR loss, where an additional regularization loss is introduced to prevent knowledge forgetting. LWF regularizes the editing process by using the MSE loss between the scores of explicit edit pairs from the original and edited models. Conversely, L2 directly employs the MSE loss between the parameters of the original model and the edited model for regularization. The formulations of their edit losses are as follows:
(8) $\mathcal{L}=\mathcal{L}_{ebpr}+\lambda\mathcal{L}_{reg}.$
@@
SRIU: SRIU normalizes the output scores of the original model on explicit edit pairs using mean and variance normalization. Then, it computes the weight for each explicit edit pair through Softmax. For the E-BPR loss generated for each triplet (u, $i_{pos}$, $i_{neg}$), the corresponding loss is multiplied by the weight of its associated explicit edit pair.
@@
RSR and SPMF: RSR selects $N$ sets of positive feedback at random from historical training data. Conversely, SPMF calculates weights based on the output scores of the original model on the training data and performs weighted sampling to select $N$ sets of positive and negative feedback. Editing is then conducted using the BCE loss.
@@
SML: The original SML combines the parameters of the original model with those trained on incremental data as inputs for the parameter transformer. However, in our task, due to the small number of explicit edit pairs, which only contain negative feedback, it is impractical to train a reasonable recommendation model. Therefore, we modify the parameters trained on incremental data to be those obtained from fine-tuning the original model on explicit edit pairs for one round.
SiReN : SiReN requires an additional MLP module to process negative feedback. After obtaining the negative feedback embeddings, they are merged with the positive feedback embeddings using an attention mechanism. The MLP module and attention parameters are pre-trained by simulating real negative feedback through random sampling of negative feedback instances.
@@
SiGRec: SiGRec aggregates negative embeddings through messages passing on a negative graph, which are then concatenated with the user and item embeddings obtained from the original collaborative filtering model. This process results in new user and item embeddings. SiGRec uses the Sign Cosine loss function proposed in the original paper for editing
@@
EGNN and BiEGNN: EGNN and BiEGNN employ an MLP as the editing module, with the user and item embeddings obtained from the original collaborative filtering model serving as inputs to the editing module. Before editing, it is necessary to pre-train the editing module. This pre-training is conducted by minimizing the score differences between the original model and the modified model on the training data, using the MSE loss.
A.4. Editing Method Parameter Setting
@@
A.5. Comparison of different recommendation editing methods.
Table 6 show the difference. We conducted a comparison of various editing methods from four different dimensions, as illustrated in Table 6. “Model Agnostic” refers to the applicability of the method to models with unknown architectures, “Training Data Agnostic” indicates its suitability for models with unknown training data, and “Extra Storage” denotes whether additional storage space is required.
A.6. Hardware and software configuration
All experiments are executed on a server with 755GB main memory, and two AMD EPYC 7763 CPUs. All experiments are done with a single NVIDIA RTX 4090 (24GB). The software and package version is specified in Table 4:
@@ A.6. Hardware and software configuration
Table 4. Package configurations of our experiments.
@@
Table 5. Recommendation models performance before editing.
Table 6. Comparison of different recommendation editing methods.
Category Editing Method Editing Module Model Agnostic Training Data Agnostic Extra Storage
Finetuning-based FT Model Parameter
EFT Model Output
Regularization-based LWF Model Parameter
L2 Model Parameter
SRIU Model Parameter
Replay-based RSR Model Parameter
SPMF Model Parameter
Optimization-based SML Model Parameter
CF with Negative Feedback SiReN Model Parameter
SiGRec Model Parameter
Model Editing EGNN Additional Architecture
BiEGNN Additional Architecture
Appendix B More Experimental Results
B.1. Recommendation Models Performance Before Editing
B.2. Editing Effectiveness
Table 3 displays a detailed comparison of various methods across different datasets and backbones, showcasing the EC, EP, Recall, and NDCG metrics, we report the average value and standard deviation of 10 runs. We can find that the results are consistent with the main analysis in Section 5.2.
B.3. Editing Efficiency
Appendix C More discussion on Recommendation Editing
C.1. Necessity of Recommendation Editing in practical applications
In practical industrial recommendation systems, the focus is more on long-term user engagement (such as user retention) rather than just whether a user clicks on more content in the short term(Zou et al., 2019). The purpose of Recommendation Editing is to improve user experience by correcting inappropriate items in recommendation results, especially discriminatory and illegal ones, thereby avoiding damage to the long-term benefits of the recommendation system. Consider, for instance, if content that severely contradicts a user’s religious beliefs is recommended, it can significantly alienate the user. Upon receiving user feedback, we need to quickly correct our recommendations to prevent such inappropriate content from appearing in future browsing sessions. Otherwise, if a user repeatedly encounters inappropriate content, they are likely to discontinue interacting with the platform and may even choose to stop using it altogether.
C.2. Application of Recommendation Editing in Industrial Recommendation Systems
Recommendation Editing can effectively function as a policy and be integrated into the reranking phase of multi-stage recommendation systems, which are commonly used in the industry. The rerank stage determines the item list that the recommendation system ultimately presents to the user. In the rerank stage, multiple policies (such as diversity) are often considered comprehensively to adjust the RankScore generated in the final rank stage(Huang et al., 2021):
$RS_{rerank}=RS*\prod_{i}(\alpha_{i}*PolicyScore_{i})^{\beta_{i}},$
where $RS_{rerank}$ is the score that determines the final recommendation result, $RS$ is the score generated in the final rank stage, $PolicyScore_{i}$ is the correction score provided by the i-th policy, , $\alpha_{i}$,$\beta_{i}$ is the hyperparameters that control the weight of the i-th policy. The recommended editing task can be easily encapsulated into a policy to adjust the final RankingScore. A simple way is to normalize the rankings of each item after editing and use them as PolicyScore:
$PolicyScore_{edit}=\frac{candidateNum}{candidateNum+rank_{edit}},$
where $candidateNum$ is the size of candidate items in the rerank stage, and $rank_{edit}$ is the ranking after editing.
C.3. Further discussion on limitation of Recommendation Editing in practical applications
While our work on recommendation editing introduces innovative strategies for enhancing recommendation systems, it is crucial to recognize several limitations and potential negative impacts associated with this approach.
First, the risk of editing attacks poses a significant concern. Malicious actors could exploit the recommendation editing process by introducing false negative signals, tricking the system into misclassifying relevant items as undesirable. This could lead to the unjust removal of valuable content from recommendation lists, thereby degrading the overall quality of recommendations and undermining user trust in the system.
Second, the potential for fairness issues must be considered. The ability to edit recommendations raises the possibility of biased outcomes, where certain items or content types are unfairly favored over others. If not carefully managed, this could result in the marginalization of diverse perspectives and content, leading to a less inclusive recommendation environment. Ensuring that recommendation editing processes are transparent and equitable is essential to mitigate this risk.
Finally, the phenomenon of catastrophic forgetting poses a significant risk. In dynamically evolving environments, continual edits to the recommendation system could inadvertently cause the system to forget previously learned preferences or associations. This could diminish the system’s ability to provide consistent and reliable recommendations over time, especially in cases where long-term user preferences are critical.